OpenStudy (anonymous):

Write an equation in slope-intercept form for an equation of a line perpendicular to y = 24x +2 and that intersects at point (0, -2). Answer y = -1/24x + 2 y = 1/24x + 2 y = -1/24x – 2 y = 1/24x – 2

OpenStudy (anonymous):

@privetek

OpenStudy (anonymous):

perpendicular means inverse of the slope so m=-24 y-y1=m(x-x1) y-(-2)=-24(x-0) y+2=-24x y=-24x-2

OpenStudy (anonymous):

@privetek, you're too fast for me

OpenStudy (anonymous):

sorry.. -1/24